# A Multi-Scale Feature Fusion Framework Integrating Frequency Domain and Cross-View Attention for Dual-View X-ray Security Inspections

### v2

@@ Abstract
With the rapid development of modern transportation systems and the exponential growth of logistics volumes, intelligent X-ray-based security inspection systems play a crucial role in public safety. Although single-view X-ray equipment is widely deployed, it struggles to accurately identify contraband in complex stacking scenarios due to strong viewpoint dependency and inadequate feature representation. To address this, we propose an innovative multi-scale interactive feature fusion framework tailored for dual-view X-ray security inspection image classification. The framework comprises three core modules: the Frequency Domain Interaction Module (FDIM) enhances frequency-domain features through Fourier transform; the Multi-Scale Cross-View Feature Enhancement (MSCFE) leverages cross-view attention mechanisms to strengthen feature interactions; and the Convolutional Attention Fusion Module (CAFM) efficiently fuses features by integrating channel attention with depthwise-separable convolutions. Experimental results demonstrate that our method outperforms existing state-of-the-art approaches across multiple backbone architectures, particularly excelling in complex scenarios with occlusions and object stacking. The code is available in anonymous repositories: https://anonymous.4open.science/r/ijcnn2025-C56D/.
@@ I Introduction
As modern transportation systems develop rapidly and logistics volumes grow exponentially, public security faces increasingly severe challenges [1]. In this context, X-ray security inspection systems play a crucial role in detecting prohibited items and ensuring public safety. X-ray inspection equipment is widely used in airports, train stations, customs, post offices, and other public places to examine luggage, parcels, and other items, ensuring they do not contain dangerous goods, contraband, or explosives. However, with the growing complexity of transportation systems, traditional single-view X-ray inspection equipment has become inadequate to meet the increasing security demands.
The widespread use of single-view X-ray inspection equipment has provided basic security guarantees but still faces inherent problems. Due to the limitations of single-view imaging principles, X-rays pass through packages, and materials of varying compositions, densities, and thicknesses absorb X-rays differently, affecting the display of images. Especially in complex stacking scenarios, single-view equipment may not clearly present the complete information of target objects, thereby affecting the detection effectiveness of prohibited items [2, 3, 4, 5, 6].
With technological advancements, dual-view X-ray inspection technology has gradually become an important means to address these problems. By simultaneously imaging objects from two orthogonal directions, dual-view X-ray inspection technology significantly enhances the spatial representation of target objects. This is particularly beneficial in complex stacking, angular deviations, and object occlusions, as it can provide more complete contours and morphological information of prohibited items, thereby improving detection accuracy [7]. However, despite significant progress in the practical application of dual-view X-ray equipment, most commercial systems still adopt independent dual-channel detection modes and fail to fully exploit the complementary features between the two views. This results in cross-modal collaborative detection performance far below its theoretical optimum [8].
Figure 1: This image demonstrates the advantages of dual-view, as it reveals features that are not visible in a single view.
@@
To address these technical bottlenecks, this paper proposes a multi-scale interactive feature fusion framework for dual-view X-ray security inspection images. The framework comprises three innovative core modules:
  • Frequency Domain Interaction Module (FDIM): Extracts frequency-domain features via the Fourier transform and establishes cross-modality feature associations in the frequency domain, effectively enhancing the complementary expression of fine-grained details.
  • Multi-Scale Cross-View Feature Enhancement (MSCFE): Employs cross-view attention networks to adaptively strengthen feature interactions between views, improving target discrimination in complex scenarios.
  • Convolutional Attention Fusion Module (CAFM): Incorporates channel attention mechanisms and depthwise-separable convolution units to precisely fuse dual-view features while suppressing redundancy.
Through systematic experiments on mainstream backbones, ablation studies demonstrate that each module significantly boosts performance. Comparative experiments further show that our model outperforms state-of-the-art methods across multiple backbones.
@@ II Related Research
In the field of X-ray security image processing, most existing approaches are designed from a single-view perspective, predominantly enhancing feature representations to tackle challenges such as occlusion and complex backgrounds. For instance, the Occlusion-Aware Attention Module (DOAM) [9] focuses on extracting edge information from images, reducing the interference caused by occluded objects by concentrating on their edge features. CFPA-Net [10] employs cross-layer feature fusion (CEF) to reinforce the interaction between high-level semantic features and low-level location cues, thereby improving detection accuracy in complex backgrounds. IEFPN [11] introduces a Layer-based Recalibration Module (LRM) to optimize multi-scale feature fusion on top of a traditional Feature Pyramid Network (FPN), resulting in better detection of targets with varying scales. Methods such as Re-BiFPN [12], and FDTNet [13] have also proposed innovations in feature pyramids, attention mechanisms, and dual-stream networks (combining frequency and spatial domains), respectively, and have seen notable performance gains. However, because these methods are all single-view solutions, they are still constrained by insufficient viewpoint information and struggle with large-angle occlusions or overlapping objects in more complex scenarios.
@@
To further overcome the limitations of single-view methods, researchers have begun exploring dual-view data and networks. With the introduction of dual-view datasets [14], the AHCR [7] approach leverages information exchange between two viewpoints, achieving better detection accuracy in occluded and complex environments. Nevertheless, the existing methodology does not fully exploit the unique characteristics of both dual-view and X-ray security images, leaving room for improvement in situations with extreme occlusions or multiple overlapping items. Some researchers have proposed Transformer-based dual-view methods [15] by refining the attention mechanism for inter-view information exchange. However, these methods exhibit weak generalizability, making them difficult to adapt across various model architectures.
@@
Figure 3: The sub-modules of the proposed framework: (a) Frequency Domain Interaction Module, (b) Multi-Scale Cross-View Feature Enhancement, (c) Convolutional Attention Fusion Module.
@@ III-A Overall Architecture
As illustrated in Figure 2, the proposed dual-view X-ray image processing framework is designed to maximize feature extraction efficiency while minimizing computational overhead. We analyze the channel and resolution variations of the backbone network and divide it into four primary modules (some backbones, such as ResNet, may include five modules if a stem layer is counted). Considering the partial similarity between the two views and the need to save computational resources, we employ a weight-sharing backbone structure.
Right after the first module, we introduce the Frequency Domain Interaction Module (FDIM), which extracts frequency-domain features through the Fourier transform and establishes cross-modality feature associations in the frequency domain. This process mainly operates at a relatively higher resolution to enhance fine-grained details. It effectively compensates for the shortcomings of traditional spatial methods in detail capture, thus providing richer feature representations for subsequent fusion and detection.
Subsequently, each of the following modules adopts the Multi-Scale Cross-View Feature Enhancement (MSCFE) mechanism to perform feature interaction. In this module, we utilize cross-view attention to enhance feature interactions between different views and introduce adaptive multi-scale feature enhancement to further boost the model’s capability in discriminating targets under complex scenarios. The cross-view attention mechanism dynamically adjusts the depth of feature interactions, enabling the model to capture key semantic information at multiple scales while effectively addressing differences in viewpoint.
Finally, we design the Convolutional Attention Fusion Module (CAFM) as the last step for refined feature fusion. By leveraging channel attention mechanisms and depthwise-separable convolution units, this module meticulously fuses dual-view features while suppressing redundancy. Through this optimization, the final output has both high discriminative power and avoids superfluous information. With carefully designed module interactions and feature enhancement mechanisms, our framework can maintain computational efficiency while improving the classification performance of dual-view X-ray images. The collaboration among these modules enables the model to provide more accurate detection and classification results, particularly in scenarios with heavily stacked objects and occlusions.
In the following sections, we will detail the design of each module.
@@ III-B Frequency Domain Interaction Module (FDIM)
In dual-view X-ray image processing, discrepancies in viewpoints and overlapping objects often lead to feature misalignment and information loss. FDIM is designed to address these challenges by extracting and enhancing features in the frequency domain, leveraging both global (low-frequency) and local (high-frequency) information for more discriminative representations. This approach is inspired by the work of Tatsunami et al. [16], who explored the effectiveness of frequency-domain enhancement to mitigate feature misalignment due to viewpoint discrepancies.
III-B1 Frequency Transform
@@
As shown in the Figure 3-(a), FDIM first applies Fast Fourier Transform (FFT) [17] to map the input feature maps $x$ and $y$ from the spatial domain into the frequency domain:
$\mathit{\tilde{x}}=\mathcal{F}(\mathit{x}),\quad\mathit{\tilde{y}}=\mathcal{F}(\mathit{y})$ (1)
where $\mathcal{F}(\cdot)$ denotes the 2D FFT.
III-B2 Dynamic Filtering Enhancement
To enhance discriminative information, FDIM applies dynamic filtering using learnable weight matrices $W_{x}$ and $W_{y}$:
$\mathit{\tilde{x}}^{\prime}=\mathit{\tilde{x}}\cdot W_{x},\quad\mathit{\tilde{y}}^{\prime}=\mathit{\tilde{y}}\cdot W_{y}$ (2)
where $W_{x}$ and $W_{y}$ are dynamically computed via an MLP-based weighting mechanism:
$r_{x}=\mathit{P_{m}(x)},\quad r_{y}=\mathit{P_{m}(y)}$ (3)
where $P_{m}(\cdot)$ denote average pooling operations.
$W_{x}=\text{Softmax}(\text{MLP}(r_{x}))$ (4)
$W_{y}=\text{Softmax}(\text{MLP}(r_{y}))$ (5)
III-B3 Inverse Transform and Feature Fusion
The enhanced frequency-domain features are then transformed back to the spatial domain and fused with the original features using a residual connection:
$x^{\prime}=\mathcal{F}^{-1}(\mathit{\tilde{x}}^{\prime})+x$ (6)
$y^{\prime}=\mathcal{F}^{-1}(\mathit{\tilde{y}}^{\prime})+y$ (7)
where $\mathcal{F}^{-1}(\cdot)$ denotes the inverse 2D FFT.
This process preserves both spatial contextual information and the fine-grained details extracted in the frequency domain. Compared to methods solely operating in either the spatial or frequency domain, FDIM enables flexible integration of shape and deformation cues across dual-view images, forming a robust foundation for subsequent multi-scale interactions and attention-based fusion.
III-C Multi-Scale Cross-View Feature Enhancement (MSCFE)
Due to viewpoint discrepancies and potential object overlaps, dual-view X-ray images (vertical view and horizontal view) often exhibit varying resolutions and complex scenes, posing significant challenges to object recognition. To effectively address these issues, the proposed MSCFE module(Figure 3-(b)) integrates cross-view attention mechanisms with multi-scale information fusion, aiming to balance global and local features while mitigating the effects of occlusions and cluttered backgrounds.
Firstly, the MSCFE module injects positional encoding into the input features of the vertical view branch $f_{ol}$ and the horizontal view branch $f_{sd}$. Specifically, a 2D sinusoidal positional encoder generates positional embeddings, which are then added to the input features to explicitly embed spatial positional information:
$F_{ol}^{P}=F_{ol}+\text{PosEnc}(F_{ol})$ (8)
$F_{sd}^{P}=F_{sd}+\text{PosEnc}(F_{sd})$ (9)
where $F_{ol}$ denotes the input features from the vertical view branch, $F_{sd}$ denotes the input features from the horizontal view branch, and $\text{PosEnc}(\cdot)$ represents the positional encoding function.
Subsequently, the MSCFE module employs a Multi-Head Cross-Attention mechanism [18, 19, 20] to facilitate feature interactions between the vertical and horizontal views. Specifically, the vertical view branch features are used as queries (Query), while the horizontal view branch features serve as keys (Key) and values (Value), and vice versa, enabling bidirectional feature enhancement:
$F_{sd}^{att}=\text{CrossAttn}(F_{ol}^{P},F_{sd}^{P})$ (10)
$F_{ol}^{att}=\text{CrossAttn}(F_{sd}^{P},F_{ol}^{P})$ (11)
where $F_{sd}^{att}$ and $F_{ol}^{att}$ are the attention-enhanced feature representations for the horizontal and vertical view branches, respectively.
Following the cross-attention operations, the enhanced features are fused with the original input features and subjected to Batch Normalization to maintain feature stability:
$F_{ol}^{E}=\text{BatchNorm}(F_{ol}^{att})+F_{ol}$ (12)
$F_{sd}^{E}=\text{BatchNorm}(F_{sd}^{att})+F_{sd}$ (13)
This approach not only preserves the semantic integrity of the original features but also incorporates useful information discovered through cross-view interactions.
Building upon the fused features, the MSCFE module further refines the representations using Depthwise Separable Convolution (DWConv [21]) to enhance edge and detail representations:
$F_{ol}^{conv}=\text{DWConv}(F_{ol}^{E})$ (14)
$F_{sd}^{conv}=\text{DWConv}(F_{sd}^{E})$ (15)
where $F_{ol}^{conv}$ and $F_{sd}^{conv}$ are the processed features for the vertical and horizontal view branches, respectively.
Additionally, when the module is not the last layer, the MSCFE module introduces optional feature downsampling layers. These layers employ Depthwise Separable Convolutions and pooling operations to construct spatial attention maps. The intermediate pooling results are computed as follows:
$s_{ol}^{pool}=\text{Concat}(P_{m}(F_{ol}^{conv}),P_{a}(F_{ol}^{conv}))$ (16)
$s_{sd}^{pool}=\text{Concat}(P_{m}(F_{sd}^{conv}),P_{a}(F_{sd}^{conv}))$ (17)
where $P_{m}(\cdot)$ and $P_{a}(\cdot)$ denote max pooling and average pooling operations.
Next, a $7\times 7$ convolution (denoted as $\text{Conv}_{7}$) is applied to extract spatial relationships:
$s_{ol}^{spatial}=\text{Conv}_{7}(s_{ol}^{pool})$ (18)
$s_{sd}^{spatial}=\text{Conv}_{7}(s_{sd}^{pool})$ (19)
Finally, activation functions such as Sigmoid are applied to generate the spatial attention maps:
$s_{ol}=\text{Sigmoid}(s_{ol}^{spatial})$ (20)
$s_{sd}=\text{Sigmoid}(s_{sd}^{spatial})$ (21)
Here, $s_{ol}$ and $s_{sd}$ are the attention maps generated for the vertical and horizontal views, respectively. These attention maps are employed in subsequent layers to weight the features, thereby further enhancing important features and suppressing irrelevant information.
Finally, the MSCFE module outputs the enhanced vertical and horizontal view features along with their corresponding attention maps $s_{ol}$ and $s_{sd}$. These attention maps are utilized in the following layers to perform weighted feature processing, resulting in more refined and discriminative feature representations.
In summary, the MSCFE module effectively enhances the feature representation capabilities of dual-view (vertical and horizontal) X-ray images by incorporating positional encoding, multi-head cross-view attention mechanisms, and depthwise separable convolutions. This design significantly improves target recognition accuracy and robustness, especially in complex scenarios involving occlusions and object stacking, while the generated attention maps further optimize subsequent feature processing.
III-D Convolutional Attention Fusion Module (CAFM)
After the Frequency Domain Interaction Module (FDIM) and the Multi-Scale Cross-View Feature Enhancement (MSCFE) module incorporate frequency-domain enhancement and cross-view attention, the high-resolution and low-resolution feature maps may still contain redundant or noisy elements. To further refine feature expressiveness while minimizing computational overhead, we introduce the CAFM, as shown in the Figure 3-(c), which combines channel attention mechanisms and depthwise-separable convolutions to achieve efficient and discriminative fusion of dual-view features.
Firstly, the CAFM module applies the Convolutional Block Attention Module (CBAM [22]) to the feature maps from different resolution branches:
$F_{ol}^{\prime}=\text{CBAM}(F_{ol})$ (22)
$F_{sd}^{\prime}=\text{CBAM}(F_{sd})$ (23)
where $F_{ol}^{\prime}$ and $F_{sd}^{\prime}$ are the attention-enhanced outputs of the vertical and horizontal view features, respectively. CBAM computes attention weights along both the channel and spatial dimensions, effectively filtering out redundant information while highlighting critical features.
Next, the CAFM module concatenates these attention-refined feature maps along the channel dimension and applies normalization:
$F_{cat}=\text{concat}(F_{ol}^{\prime},F_{sd}^{\prime})$ (24)
$F_{bn}=\text{BatchNorm}(F_{cat})$ (25)
This fusion ensures the integrity of dual-view information while stabilizing the feature distribution for subsequent convolution operations.
Finally, the CAFM module processes the concatenated and normalized features through a dwcsplayer, followed by a ReLU activation function:
$F_{csp}=\text{dwcsplayer}(F_{bn})$ (26)
$F_{final}=\text{ReLU}(F_{csp})$ (27)
where $F_{csp}$ is the feature representation processed by the dwcsplayer, and $F_{final}$ is the final output feature map used for classification or detection tasks.
The dwcsplayer integrates a single Depthwise Separable Convolution (DWConv) with feature fusion to effectively refine and enhance feature representations while maintaining computational efficiency [23]. ReLU activation effectively suppresses redundant information and retains the most discriminative features.
The dwcsplayer plays a crucial role within the Convolutional Attention Fusion Module (CAFM). It integrates a single DWConv with feature fusion to effectively enhance and refine feature representations while maintaining computational efficiency. Specifically, the dwcsplayer performs the following operations:
Firstly, a $1\times 1$ convolutional layer (Conv1) reduces the channel dimensionality to mitigate computational complexity:
@@ III-D Convolutional Attention Fusion Module (CAFM)
Simultaneously, the input feature is processed via a DWConv to extract spatial features:
$F_{x_{1}}=\text{DWConv}(F_{bn})$ (29)
The outputs of these two feature branches are then element-wise summed to merge information from both pathways:
$F_{sum}=F_{x_{1}}+F_{x_{2}}$ (30)
This design enables dwcsplayer to efficiently refine spatial and channel-wise features while preserving essential discriminative information.
The CAFM module effectively fuses dual-view feature maps by combining channel attention mechanisms with a depthwise-separable convolution. The dwcsplayer within CAFM ensures that the fused features are refined and enhanced efficiently, preserving important spatial and channel-wise information while suppressing redundant data. This design not only improves the model’s accuracy but also ensures computational feasibility during deployment.

### v4

# DAGNet: A Dual-View Attention-Guided Network for Efficient X-ray Security Inspection

@@ Abstract
With the rapid development of modern transportation systems and the exponential growth of logistics volumes, intelligent X-ray-based security inspection systems play a crucial role in public safety. Although single-view X-ray baggage scanner is widely deployed, they struggles to accurately identify contraband in complex stacking scenarios due to strong viewpoint dependency and inadequate feature representation. To address this, we propose a Dual-View Attention-Guided Network for Efficient X-ray Security Inspection (DAGNet). This study builds on a shared-weight backbone network as the foundation and constructs three key modules that work together: (1) Frequency Domain Interaction Module (FDIM) dynamically enhances features by adjusting frequency components based on inter-view relationships; (2) Dual-View Hierarchical Enhancement Module (DVHEM) employs cross-attention to align features between views and capture hierarchical associations; (3) Convolutional Guided Fusion Module (CGFM) fuses features to suppress redundancy while retaining critical discriminative information. Collectively, these modules substantially improve the performance of dual-view X-ray security inspection. Experimental results demonstrate that DAGNet outperforms existing state-of-the-art approaches across multiple backbone architectures. The code is available at: https://github.com/ShilongHong/DAGNet.
@@ I Introduction
As modern transportation systems develop rapidly and logistics volumes grow exponentially, public security faces increasingly severe challenges[1]. In this context, X-ray security inspection systems play a crucial role in detecting prohibited items and ensuring public safety. X-ray inspection equipment is widely used in airports, train stations, customs, post offices, and other public places to examine luggage, parcels, and other items, ensuring they do not contain dangerous goods, contraband, or explosives. However, due to the growing complexity of transportation systems, traditional single-view X-ray baggage scanner has become inadequate for meeting rising security demands.
The widespread use of single-view X-ray baggage scanner has provided basic security guarantees but still faces inherent problems. Due to the limitations of single-view imaging principles, X-rays pass through packages, and materials of varying compositions, densities, and thicknesses absorb X-rays differently, affecting the display of images. In complex stacking scenarios, single-view systems often fail to capture complete information about overlapping items, significantly compromising the accuracy of detecting prohibited objects.[2, 3, 4, 5, 6].
@@
With technological advancements, dual-view X-ray inspection technology has gradually become an important means to address these problems. By imaging objects simultaneously from two orthogonal directions, dual-view X-ray technology greatly improves spatial representation. This is particularly beneficial in complex stacking, angular deviations, and object occlusions, as it can provide more complete contours and morphological information of prohibited items, thereby improving detection accuracy[7]. Although these methods can alleviate the viewpoint difference problem and supplement the information to a certain extent, it is difficult to fully exploit the detail complementarity between viewpoints, which makes the information redundancy and lack of details due to the viewpoint difference, and ultimately leads to the detection effect being adversely affected.
Figure 2: Overall architecture of the proposed framework.
To address these challenges, we propose DAGNet, an attention-guided network designed specifically for dual-view X-ray security screening. We formulate the task as a multi-label, multi-category classification problem, effectively overcoming geometric variations and modality-specific differences inherent in dual-view imaging. This task definition can effectively avoid the limitations of the detection paradigm due to cross-view geometric variations and modality-specific differences, making the model more focused on attribute-level feature associations between views, and has become one of the mainstream solutions in current related research. Different from the existing approaches, the main contribution of this study lies in the design of a systematic multi-level feature interaction architecture, which effectively solves the problem of information loss in dual-view fusion.
In the design of the model, this study takes the backbone network with shared weights as the basis and constructs three mutually synergistic key modules as the main innovations:
  • The Frequency Domain Interaction Module (FDIM), dynamically enhancing features by adjusting frequency components based on inter-view relationships
  • The Dual-View Hierarchical Enhancement Module (DVHEM), employing cross-attention mechanisms to adaptively align and integrate features across views at multiple scales
  • The Convolutional Guided Fusion Module (CGFM), efficiently fusing features to suppress redundancy while preserving discriminative information.
@@ II Related Research
In the field of X-ray security image processing, most existing approaches are designed from a single-view perspective, predominantly enhancing feature representations to tackle challenges such as occlusion and complex backgrounds. For instance, the De-Occlusion Aware Attention Module (DOAM)[9] focuses on extracting edge information from images, reducing the interference caused by occluded objects by concentrating on their edge features. CFPA-Net[10] employs cross-layer feature fusion to reinforce the interaction between high-level semantic features and low-level location cues, thereby improving detection accuracy in complex backgrounds. IEFPN[11] introduces a Layer-based Recalibration Module (LRM) to optimize multi-scale feature fusion on top of a traditional Feature Pyramid Network (FPN), resulting in better detection of targets with varying scales. Methods such as Re-BiFPN[12], and FDTNet[13] have also proposed innovations in feature pyramids, attention mechanisms, and dual-stream networks (combining frequency and spatial domains), respectively, and have seen notable performance gains. However, these single-view methods are constrained by the lack of complementary viewpoint information, leading to difficulties in addressing severe occlusions and overlapping objects from extreme viewing angles.
@@
To further overcome the limitations of single-view methods, researchers have begun exploring dual-view data and networks[14]. With the introduction of dual-view datasets, the Attentive Hierarchical Cross-View Representation (AHCR)[7] approach leverages information exchange between two viewpoints, achieving better detection accuracy in occluded and complex environments. Nevertheless, current dual-view methodologies do not adequately leverage the inherent structural characteristics and complementary details of X-ray security images, especially in complex scenarios with extreme occlusions or multiple overlapping items. Some researchers have proposed Transformer-based dual-view methods[15] by refining the attention mechanism for inter-view information exchange. However, these methods exhibit weak generalizability, making them difficult to adapt across various model architectures.
@@ III-A Overall Architecture
As illustrated in Figure 2, the DAGNet proposed in this work aims to maximize feature extraction efficiency and achieve efficient and accurate security inspection tasks. The entire framework is divided into four main blocks based on changes in the size of the feature maps output by the backbone network. Considering that the feature information of the two perspectives partially overlaps and complements each other, this study adopts a backbone network structure with shared weights to ensure that dual-perspective feature extraction remains efficient while allowing the backbone network to use as much data as possible, making the model more robust and consistent. The specific process is as follows.
First, given a pair of dual-view images, two shared backbone capture features from different viewpoints. After the first backbone network module, the feature maps are input into FDIM, where Fourier transform is used to extract frequency domain features and establish inter-viewpoint frequency domain information associations in the frequency domain.
Next, the feature maps output from the backbone of different perspectives are sent to DVHEM for feature interaction between perspectives. Adaptive multi-scale feature enhancement weights are innovatively introduced to further improve the model’s ability to distinguish targets in complex scenes.
Finally, the high-level features obtained from the two perspectives are sent to CGFM for perspective feature fusion, and the fused features are sent to the classification head to complete the final output.
In the subsequent sections, we will elaborate on the detailed structure and operational principles of each proposed module.
@@ III-B Frequency Domain Interaction Module (FDIM)
Figure 3: The overall architecture of the proposed Frequency Domain Interaction Module (FDIM)
Capturing features of overlapping items is challenging in dual-view X-ray imaging due to its transmissive imaging nature. To effectively address this issue, we propose the Frequency Domain Interaction Module (FDIM), illustrated in Figure 3, leveraging frequency-domain analysis to enhance feature extraction capabilities.
Unlike previous frequency domain methods, FDIM dynamically adjusts and enhances frequency components by collaboratively constructing adaptive filters from dual-view features. Specifically, important frequency bands are amplified while noisy bands are suppressed, effectively leveraging global and frequency-selective properties to yield more discriminative features compared to spatial-domain approaches.
@@
FDIM first performs Pointwise Convolution (PWConv) [16] on the input feature map $x$ to integrate channel information, and then applies a two-dimensional fast Fourier transform[17] to map it from the spatial domain to the frequency domain:
@@
[…]
Subsequently, to determine the appropriate filter parameters, inspired by Tatsunami et al.[18] in their work on frequency domain, this design divides the filter D into two parts: a learnable base filter and a channel-wise weighted weight matrix to better enhance feature expression.
First, the filter weight matrix $W_{x}$ is generated through interactions between dual-view input features, constructing a learnable channel weighting mechanism that emphasizes key channel responses. This is formally defined as:
$x^{\prime}=P_{m}(x)\cdot ViewMix(x,y)+P_{m}(x)$ (2)
$ViewMix(x,y)=Sigmoid(PWConv(P_{m}(y)))$ (3)
[…]
Subsequently, the learned base filter $K$ is modulated at the channel level to produce the adaptive frequency-domain filter $\mathit{D_{X}}$:
$\mathit{D_{X}}=\mathit{{K}}\cdot W_{x}$ (5)
where $\mathit{D_{X}}$ is the generated frequency domain filter, $\mathit{K}$ is the learnable basic filter, a complex matrix with dimensions $H\times\lfloor W/2\rfloor$, and $W_{x}$ is the channel-level weight matrix used to adjust the response intensity of different channels in the frequency domain.
The enhanced frequency-domain feature map is then transformed back to the spatial domain using inverse FFT and integrated through channel-wise convolution. Finally, residual connections stabilize feature learning:
$F_{X}=PWConv(\mathit{F}^{-1}(\mathit{{F_{X}}}^{\prime}\cdot\mathit{D_{X}}))+x$ (6)
[…]
III-C Dual-View Hierarchical Enhancement Module (DVHEM)
In security inspection scenarios, dual-view X-ray images exhibit distinct feature distributions. Although the same object exhibits feature differences under different views, it still contains significant associated information. Additionally, due to the differing shapes and sizes of features presented in different views, there exists potential associativity between low-level and high-level features. To fully leverage these characteristics, this work proposes the DVHEM, as shown in Figure 4. The module consists of two key components: first, it enhances the feature representations of each view through interactions based on the relationships between views; second, it constructs a Hierarchical Guidance Attention module (HGA) that guides attention across levels. This design fully integrates and utilizes the correlations between features from different viewpoints and scales, balancing global information and local details, thereby significantly improving performance in dual-view security inspection scenarios.
First, DVHEM adds position encoding to the input features of the vertical viewpoint branch $F_{OL}$ and the horizontal viewpoint branch $F_{SD}$ to introduce spatial position information, as defined below:
$F_{OL}^{P}=F_{OL}+{PosEnc}(F_{OL})$ (7)
$F_{SD}^{P}=F_{SD}+{PosEnc}(F_{SD})$ (8)
Here, $F_{OL}$ and $F_{SD}$ denote the input feature maps of the vertical and horizontal view angle branches, respectively, ${PosEnc}(\cdot)$ denotes the cosine position encoding function, and $F_{OL}^{P}$ and $F_{SD}^{P}$ denote the feature maps after adding position encoding.
Figure 4: Overall architecture of the Dual-View Hierarchical Enhancement Module (DVHEM).
Next, DVHEM adopts a multi-head cross-attention mechanism[19, 20, 21] to promote effective feature interaction between vertical and horizontal viewpoints. In this mechanism, the features of one branch are used as Q (Query), and the features of the other branch are used as K (Key) and V (Value), which can be defined as:
$F_{SD}=MHA(Q=F_{OL}^{P},K=F_{SD}^{P},V=F_{SD}^{P})$ (9)
$F_{OL}=MHA(Q=F_{SD}^{P},K=F_{OL}^{P},V=F_{OL}^{P})$ (10)
where $MHA(\cdot)$ denotes the multi-head attention mechanism, $Q$ , $K$ , and $V$ are the query, key, and value in the attention mechanism, respectively.
To further enhance information transmission between different layers, the feature maps obtained from cross-attention are input into the proposed HGA module. First, this module uses Depthwise Separable Convolution (DWConv) [22, 23] for preliminary feature extraction and dimension reduction to reduce the computational burden while highlighting feature regions, as expressed below:
$F_{OL}^{down}={AvgPool}({DWConv}(F_{OL}))$ (11)
where $F_{OL}$ is the vertical view angle branch feature map, $DWConv(\cdot)$ denotes depth-separable convolution, $AvgPool(\cdot)$ denotes average pooling, which reduces the feature map resolution and enhances regional aggregation capabilities, and $F_{OL}^{down}$ denotes the processed intermediate features.
Next, channel-dimensional information extraction and enhancement are performed on the feature maps to capture a more comprehensive feature representation, as follows:
$F_{OL}^{Channel}=Concat(C_{M}(F_{OL}^{down}),C_{A}(F_{OL}^{down}))$ (12)
where $C_{M}(\cdot)$ and $C_{A}(\cdot)$ denote the maximum pooling (ChannelMax) and average pooling (ChannelAvg) operations in the channel dimension, respectively, and $Concat(\cdot)$ denotes the feature concatenation operation. The output $F_{OL}^{Channel}$ is the feature map enhanced at the channel level.
Finally, a 7×7 convolution is used to further enhance the spatial feature representation capability, and the output is mapped to the range (0, 1) through the Sigmoid activation function to generate the guidance attention weight maps $S_{OL}$ and $S_{SD}$, which are then passed to the next layer through matrix multiplication. This ensures that the model can obtain more global structural information and image details. Taking the generation of $S_{OL}$ as an example, this process can be expressed as:
$S_{OL}=Sigmoid(Conv_{7}(F_{OL}^{Channel}))$ (13)
where $Conv_{7}(\cdot)$ denotes a convolution operation with a kernel size of $7\times 7$, $Sigmoid(\cdot)$ denotes the Sigmoid activation function, and $S_{OL}$ denotes the generated guidance attention weight map.
III-D Convolutional Guided Fusion Module (CGFM)
Directly concatenating or simply adding together feature maps from different sources often introduces redundant information or noise, thereby reducing the discriminative ability of the features. To solve this problem and efficiently fuse heterogeneous features from two perspectives, this study proposes CGFM, which combines spatial–channel attention with a dual-branch bottleneck fusion to produce highly discriminative features.
As shown in Figure 5, this module introduces a Convolutional Block Attention Module (CBAM[24]) and Depthwise Separable Convolution (DWConv) to achieve dual-view feature fusion in an efficient and discriminative manner, thereby enhancing the detection capability of the model.
Formally, let $F_{OL}$ and $F_{SD}$ represent the feature maps of the two views. First, the CBAM processes the two feature maps separately, adaptively adjusting the importance of different regions and channels to highlight key features and suppress redundant information:
$F_{OL}^{\prime}=CBAM(F_{OL}),\quad F_{SD}^{\prime}=CBAM(F_{SD})$ (14)
where $F_{OL}$ and $F_{SD}$ represent the original feature maps from vertical and horizontal views, respectively, and $CBAM(\cdot)$ denotes the channel-spatial attention mechanism, which enhances discriminative regional features while suppressing redundant or noisy features, outputting $F_{OL}^{\prime}$ and $F_{SD}^{\prime}$ as attention-optimized feature maps.
Subsequently, the optimized feature maps are concatenated along the channel dimension and normalized to generate the preliminary fused feature map $F_{bn}$:
$F_{bn}=BatchNorm(Concat(F_{OL}^{\prime},F_{SD}^{\prime}))$ (15)
where $BatchNorm(\cdot)$ denotes batch normalization, $Concat(\cdot)$ denotes concatenation of attention-enhanced features from two views along the channel dimension, and $F_{bn}$ denotes the normalized fusion feature map.
Next, following the classic bottleneck structure design, the features are sent to two branches for processing: one uses DWConv to retain spatial structural information, and the other uses 1×1 convolution to integrate cross-channel information. Finally, the results of the two branches are summed. This design reduces complexity while balancing the expression of spatial features and channel relationships. The specific process can be represented as follows:
$F_{out}=DWConv(F_{bn})+Conv_{1}(F_{bn})$ (16)
where $DWConv(\cdot)$ denotes the depth-separable convolution operation, $Conv_{1}(\cdot)$ denotes the $1\times 1$ convolution operation, and $F_{out}$ denotes the final output feature map after fusion.
Figure 5: The overall architecture of the proposed Convolutional Attention Fusion Module (CGFM)